### v2

# Generative Human Geometry Distribution

Abstract
Realistic human geometry generation is an important yet challenging task, requiring both the preservation of fine clothing details and the accurate modeling of clothing-pose interactions. Geometry distributions, which can model the geometry of a single human as a distribution, provide a promising representation for high-fidelity synthesis. However, applying geometry distributions for human generation requires learning a dataset-level distribution over numerous individual geometry distributions. To address the resulting challenges, we propose a novel 3D human generative framework that, for the first time, models the distribution of human geometry distributions. Our framework operates in two stages: first, generating the human geometry distribution, and second, synthesizing high-fidelity humans by sampling from this distribution. We validate our method on two tasks: pose-conditioned 3D human generation and single-view-based novel pose generation. Experimental results demonstrate that our approach achieves the best quantitative results in terms of realism and geometric fidelity, outperforming state-of-the-art generative methods.
Figure 1: (a) Our method synthesizes high-fidelity human geometry using a diffusion denoising model. Building on this model, we propose a versatile generative framework and perform two generative tasks. (b) Pose-conditioned random generation: given a human pose, we randomly synthesize realistic geometry with natural clothing details. (c) Novel pose generation: given a normal image, we synthesize high-quality geometry in novel poses. Our results showcase detailed and pose-aware clothing deformations. All visual results are rendered from point clouds.
1 Introduction
3D human geometry generation is a critical yet challenging task. The human body exhibits high-frequency clothing details, which are inherently difficult to synthesize. Therefore, the core of this task is the design of a representation that can capture both underlying structure and fine details, posing two key challenges: 1) encoding high-frequency details into a low-dimensional manifold without losing fidelity; 2) modeling the relationship between clothing wrinkles and poses to preserve realistic details.
Existing methods represent the human body in different ways. NeRFs [23, 53, 38, 14, 45, 58] focus on the rendering results but neglect underlying geometry and are constrained by the rendering speed and resolution. Implicit functions [41], such as signed distance functions struggle to synthesize thin structures and tend to oversmooth results. Point clouds [55, 56] and volume-based representations trade-off between memory efficiency and quality. Recently, geometry distributions [52] model each geometry as a distribution, enabling “infinite” sampling from the distribution to construct high-fidelity geometry, offering a promising approach for detailed human representation. However, extending a single-geometry distribution, trained on a single shape, to an entire dataset is non-trivial and computationally inefficient, as it requires modeling the dataset as a distribution over numerous geometry distributions, as illustrated in Fig. 2, introducing hierarchical complexity.
Figure 2: The dataset distribution represents the distribution of geometries within the dataset, where each sample corresponds to a geometry. The geometry distribution illustrates the distribution of points within a single geometry. Most generative methods model the dataset distribution only, while our method models the dataset distribution of geometry distributions.
To this end, we propose the first 3D generative method that models the distribution of individual human geometry distributions. Specifically, we encode each human geometry distribution into a feature map, formulating 3D human generation as a two-stage process: first, the human geometry distribution (i.e., the feature map) is generated, and then, a high-fidelity human is synthesized by sampling from this distribution. Unlike the prior geometry distributions [52], which learn transformations from Gaussian noise to a selected shape, our key insight is to leverage the SMPL model distribution, which is inherently closer to human geometry. Instead of randomly sampling points from the SMPL model for training, we construct a mapping based on the proximity between points on the SMPL model and the human geometry surface, avoiding the inefficiencies of learning transformations between distant points. Furthermore, based on this mapping, we normalize both the SMPL and human distributions to reduce modeling complexity further. Given the distribution formulation, an auto-decoder encodes each geometry into a compact feature map, which, when combined with a human pose, represents a fine-grained human geometry capable of capturing realistic clothing details that naturally conform to body articulation.
Our method synthesizes high-fidelity geometry directly, without relying on techniques like normal maps to enhance rendering results. We design two generative tasks to evaluate the effectiveness of our method. The first task generates random 3D human geometries conditioned on a given pose, while the second enables novel pose generation from a single-view normal image. Our quantitative results show that our method improves the geometry quality by $57\%$ ($42.9$ to $16.2$) and the visual appearance by $7\%$ ($17.4$ to $16.2$) compared to the state-of-the-art. Our contribution can be summarized as follows:
  • We propose the first 3D generative method that models the distribution of individual human geometry distributions, synthesizing high-fidelity avatars with realistic clothing details, including loose parts.
  • We introduce a novel framework for modeling human geometry distributions that samples from the SMPL model, significantly reducing modeling complexity and improving training efficiency.
  • We demonstrate the effectiveness of our geometry distribution in downstream tasks, achieving superior performance compared to existing generative methods.
2 Related work
2.1 General 3D Generation
3D generation methods employ various representations, such as meshes, voxels, point clouds, and implicit functions, each offering distinct advantages but also facing inherent limitations [52]. Here, we focus on compact representations that scale efficiently to large datasets.
Given one or more images, direct inference of pixel-aligned positions [48, 32, 43, 42] avoids intermediate representations, which is unsuitable for modeling intrinsic human shapes. A popular alternative is to synthesize tri-planes as intermediate representations [33, 19, 12], from which implicit functions or points [60, 9] can be sampled. However, tri-planes are limited by their resolution, restricting their ability to capture fine-grained details. GaussianCube [51] transfers Gaussians into a regular voxel grid via optimal transport, limiting its flexibility for intricate articulation. Gamba [31] and VectorSet [50] employ cross-attention mechanisms to inject 3D embeddings or image tokens, albeit constrained by memory consumption. Trellis [37] leverages sparse voxels but lacks mechanisms to incorporate human-specific priors. In contrast, 2D maps [25, 7, 46], offer a memory-efficient alternative that aligns naturally with human priors.
2.2 3D Human Reconstruction and Generation
Features extracted from images [26, 27, 2] provide rich view-dependent information, enabling high-fidelity human reconstruction. Some works incorporate additional cues such as voxel grids [10], human templates [59], or multi-view images [30] to mitigate the depth ambiguity of single-view image. For example, ICON [39] and ECON [40] predict the frontal and back normals from the SMPL template and reconstruct the human using these maps. Building on these reconstruction techniques, generative networks—such as GANs [38, 14] and diffusion models [29, 30, 54]—have been employed to synthesize pixel-aligned features or generate multi-view images [18] for human generation. However, these methods are inherently limited by their reliance on view-dependent features, sensitivity to multi-view inconsistencies, and often require post-processing to fuse information from different views [54], failing to capture the intrinsic 3D shape.
To learn intrinsic 3D representations, a popular approach [24, 53, 47, 35, 36] is to train a GAN model that synthesizes tri-planes and uses NeRF for rendering. However, these methods are constrained by rendering speed and resolution, often requiring super-resolution modules [3, 6], multi-part structures [45], or refinement stages [23, 58] to achieve high-fidelity results. Rather than employing tri-planes, alternative methods incorporate human templates, alleviating sparse spatial occupancy, but struggle with loose clothing [11] or still rely on super-resolution modules [13]. Explicit representations, such as primitive volumes [5], provide efficient rendering but face challenges in extracting detailed geometry. Mesh-based methods, using displacement [28] or layered surface volumes [44], are limited by fidelity and struggle with loose clothing. Point-based representations [1, 56, 55] offer flexibility for modeling thin structures but are constrained by point density. Furthermore, it can be difficult to produce accurate 3D geometry under indirect supervision from rendering. In contrast, gDNA [41] proposes learning implicit functions directly from 3D data, but synthesizing high-fidelity geometry remains challenging due to inherent representation limitations [52]. Our approach directly learns from 3D data while enabling “infinite” sampling, leading to high-fidelity human geometry synthesis. See Tab. 1.
Method Loose Clothing Scalability Fine Details
Tri-plane + NeRF [24, 53]
Template-based [11, 13]
Primitive Volumes [5]
Mesh-based [28, 44, 8]
Point-based [1, 56]
Implicit Function [41, 57]
Geometry Distributions (Ours)
Table 1: Comparison of 3D human representation methods.
3 Preliminaries
3.1 Flow Matching
Flow matching [20] is a variant of diffusion models. It constructs a flow that transforms samples from a source distribution $\mathbf{p}$ to a target distribution $\mathbf{q}$. Specifically, the training objective loss is defined as follow:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{x}_{0}\sim\mathbf{p},\mathbf{x}_{1}\sim\mathbf{q},t\in[0,1]}\left\|u_{\theta}(\mathbf{x}_{t},t)-(\mathbf{x}_{1}-\mathbf{x}_{0})\right\|$ (1)
where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+\mathbf{x}_{1}$. After training, an ordinary differential equation is solved to transition from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$, allowing us to sample $\mathbf{x}_{1}\sim\mathbf{q}$ by first sampling $\mathbf{x}_{0}\sim\mathbf{p}$.
Based on the flow matching formulation, the core of our modeling is to identify a suitable source distribution and a target distribution that aligns well with our problem. For brevity, we omit $t\in[0,1]$ in subsequent equations.
3.2 Geometry Distributions
Geometry distributions model a surface $\mathcal{M}\subset\mathbb{R}^{3}$ as a probability distribution $\Phi_{\mathcal{M}}$, where any sample $\mathbf{x}_{1}\sim\Phi_{\mathcal{M}}$ corresponds to a point on the surface. To achieve this, Zhang et al. [52] adapt a diffusion model to learn the transformation from a Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{1})$ to the target distribution of surface points. A point $\mathbf{x}_{1}\in\mathcal{M}$ on the target surface can then be obtained by solving an ODE with an initial point $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
4 Method
4.1 Overview
We model human geometry as a distribution based on the SMPL model and encode it into a compact feature map, which enables efficient generation within the dataset distribution. In this section, we first introduce our human geometry distribution formulation by identifying suitable source and target distributions. Next, we describe encoding this distribution into a feature map using an auto-decoder architecture, as depicted in Fig. 11 (a). Finally, we present generative tasks aimed at synthesizing the feature map, guided by the SMPL template, optionally incorporating additional conditioning inputs, as illustrated in Fig. 11 (b).
4.2 Human Geometry Distribution
Our insight is to replace the Gaussian distribution $\mathcal{N}(0,1)$ in the prior geometry distributions [52] with the SMPL template shape distribution $\Phi_{\mathcal{S}}$, aligning the source distribution closer to the target geometry distribution $\Phi_{\mathcal{T}}$. A naive approach can be formulated as:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{x}_{0}\sim\Phi_{\mathcal{S}},\mathbf{x}_{1}\sim\Phi_{\mathcal{T}}}\left\|u_{\theta}(\mathbf{x}_{t},t)-(\mathbf{x}_{1}-\mathbf{x}_{0})\right\|,$ (2)
where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+\mathbf{x}_{1}$. Building upon this, we propose two strategies to reduce the modeling complexity and improve training efficiency, including explicitly constructing training pairs and normalizing samples from distributions.
Training Pairs Construction.
Since the learned probability flow by Eq. 2 approximates a conditional optimal transport [21], one strategy for improving training efficiency is avoiding learning extraneous paths between distant points. To achieve this, we prioritize short flow by constructing training pairs set $\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}$ where $\mathbf{x}^{\prime}_{0}\sim\Phi_{\mathcal{S}}$ is geometrically close to $\mathbf{x}_{1}\sim\Phi_{\mathcal{T}}$. Specifically, we first sample a set of points $\{\mathbf{x}_{0}\}_{\mathcal{S}}$ on the SMPL template and a set of points $\{\mathbf{x}_{1}\}_{\mathcal{T}}$ on the target geometry. Then, for each $\mathbf{x}_{1}\in\{\mathbf{x}_{1}\}_{\mathcal{T}}$, the corresponding $\mathbf{x}_{0}^{\prime}$ is obtained by:
$\mathbf{x}^{\prime}_{0}=\arg\min_{\mathbf{x}_{0}\in\{\mathbf{x}_{0}\}_{\mathcal{S}}}||\mathbf{x}_{1}-\mathbf{x}_{0}||_{2}.$
Notably, because multiple points $\mathbf{x}_{1}$ may share the same nearest SMPL points $\mathbf{x}_{0}^{\prime}$, especially in loose or wrinkled regions, directly learning a path from $\mathbf{x}_{0}^{\prime}$ to $\mathbf{x}_{1}$ leads to under-sampled geometry. Therefore, we add a perturbation drawn from $\mathcal{N}(0,\boldsymbol{\sigma})$ to $\mathbf{x}_{0}^{\prime}$, introducing randomness to enhance sample diversity. In this formulation, the source distribution becomes $\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma})$ and the target distribution is the human geometry. Sampling from the constructed training pairs, the flow matching optimization objective is denoted by:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma}),(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})\in\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}}\left\|\cdot\right\|$ (3)
Distribution Normalization.
Learning the transformation from the SMPL template to the human geometry of a dataset yields a spatial imbalance of sampling points, which arises from two factors. First, regions like the torso exhibit high point density because they remain in similar locations across different poses, while other regions are more dynamic and thus sparsely sampled. Second, the points are distributed on 2D subspaces within 3D space, leading to inherently sparse coverage. This imbalance leads to inefficient training as the model underfits low-density regions. To address this, we normalize both the source and target distribution by subtracting $\mathbf{x}_{0}^{\prime}$. The source distribution becomes a zero-centered Gaussian distribution $\mathcal{N}(0,1)$, with $\boldsymbol{\sigma}$ set to $1$, and the target distribution is modeled as a regularized dense displacement field $\Delta\mathbf{x}=\mathbf{x}_{1}-\mathbf{x}^{\prime}_{0}$. Note that this subtraction removes the positional information of $\mathbf{x}^{\prime}_{0}$. To retain this information while avoiding the imbalance impact, we reintroduce $\mathbf{x}_{0}^{\prime}$ as a conditional signal that scales the network hidden feature, indirectly influencing the feature representation. By building flow between regularized dense spaces, we reduce the modeling complexity and achieve improved training efficiency, as demonstrated in Sec 5.1.2. The optimization objective is denoted by:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}\left\|u_{\theta}(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime})-(\Delta\mathbf{x}-\mathbf{n})\right\|,\vspace{-0.5em}$ (4)
where $\mathbf{n}\sim\mathcal{N}(0,1)$, $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})\in\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}$ and $\mathbf{x}_{t}=(1-t)\mathbf{n}+t\Delta\mathbf{x}$.
Figure 3: Overview of our method. (a) We encode a geometry into a feature map, which is decompressed with a SMPL vertex map. The decompressed representation serves as a condition for our denoising network. (b) The human generation task is formulated as the conditional generation of feature maps, guided by the SMPL vertex map, optionally incorporating additional conditioning inputs.
4.3 Conditional Distribution Encoding
The dataset for our task consists of a set of pairs $\mathcal{D}=\{(\mathcal{S},\mathcal{T})\}$, each associated with a latent representation $\mathbf{z}_{\mathcal{T}|\mathcal{S}}$, which is used to recover $\mathcal{T}$ by conditioning on $\mathcal{S}$. This design aligns with our generative task, i.e., $\mathcal{S}$-conditioned $\mathcal{T}$ generation. The training loss can be written as,
$\operatorname*{arg\,min}_{\theta,\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}}$ $\mathbb{E}_{(\mathcal{S},\mathcal{T})\in\mathcal{D}}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}$ (5)
$\left\|u_{\theta}(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime},\mathbf{z}_{\mathcal{T}|\mathcal{S}})-(\Delta\mathbf{x}-\mathbf{n})\right\|,$
To enhance the expressiveness, we model $\mathbf{z}_{\mathcal{T}|\mathcal{S}}\in\mathbb{R}^{C\times H\times W}$ as a compressed 2D feature map (height $H$ and width $W$), as shown in Fig. 11 (a). The representation aligns with recent advancements in UV representations for 3D objects [46]. Furthermore, we use a decoder network $\mathrm{Dec}_{\phi}(\cdot)$ to decompress $\mathbf{z}_{\mathcal{T}|\mathcal{S}}$ to a higher resolution. By leveraging the UV coordinates associated with $\mathbf{x}_{0}^{\prime}$, we can then obtain per-point latents on the surface $\mathcal{S}$ by sampling the decompressed map, denoted as $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Instead of directly using $\mathrm{z}_{\mathcal{T}|\mathcal{S}}$ as input, we adopt the local geometric feature $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Thus, the revised optimization can be written as
$\operatorname*{arg\,min}_{\theta,\phi,\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}}$ $\mathbb{E}_{(\mathcal{S},\mathcal{T})\in\mathcal{D}}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}$ (6)
$\left\|u_{\theta}\left(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime},\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})\right)-(\Delta\mathbf{x}-\mathbf{n})\right\|.$
Decoder.
As illustrated in Fig. 11 (a), the network $\mathrm{Dec}_{\phi}(\cdot)$ is a UNet-style network that contains several downsampling and upsampling layers. We render SMPL vertex positions into UV maps and concatenate them with hidden features in the convolution layers.
Denoising network.
The design of the denoiser $u_{\theta}$ is adapted from [52]. In this case, we have two additional conditioning signals, SMPL point $\mathbf{x}^{\prime}_{0}$ and latent $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Additionally, we augment $\mathbf{x}^{\prime}_{0}$ by concatenating normals and canonical coordinates. Surface normals provide directional cues for clothing inference, while canonical coordinates encode body part semantics (e.g., distinguishing limbs from the torso).
4.4 Generative Tasks
When we finish the learning of the latents $\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}$, we train generative models in the latent space. We investigate two tasks: pose-conditioned generation and novel pose generation from a single-view normal image, both built on a U-Net backbone [15, 16] architecture, as illustrated in Fig. 11 (b).
Pose-conditioned generation synthesizes diverse human geometries conditioned on an SMPL template mesh. To encode pose information, we render SMPL vertex positions into UV maps and inject them as residual connections into the U-Net architecture. For novel pose generation, the model additionally takes a frontal normal image as an extra condition alongside the template mesh. Specifically, for each animation sequence in the dataset, we randomly select one frame to provide the normal image condition and another frame to provide the SMPL pose condition. We leverage the frozen DINO-ViT model [4] to extract image features, which are then fused into the U-Net via cross-attention layers.
Notably, while we explicitly condition both tasks on human pose to model the inherent clothing deformation with body articulation, our model retains the flexibility to synthesize plausible geometries even when pose information is mismatched to the feature map, as demonstrated in the “Pose Variation” case of Fig. 8.
The pose-conditioned generation networks are trained on the THuman2 dataset [49], while the novel pose generation is trained on 4DDress [33]. See the Appendix for implementation details.
5 Experiments and Results
In this section, we first conduct ablation studies on the training pair construction strategy, different formulations of geometry distribution, and network architecture. Then we compare our method with state-of-the-art generative methods regarding the pose-conditioned generation. Finally, we showcase our synthesizing results of novel pose generation.
5.1 Ablation Study
5.1.1 Training Pairs Construction
Rather than searching for the nearest points on the dense SMPL mesh, we first sample a relatively sparse set of points from the SMPL mesh and then determine the nearest points $\mathbf{x}_{0}’$ ’ from this sampled set. It distributes the mapping workload across different $\mathbf{x}_{0}’$ ’, which improves sampling efficiency. To validate this, we train our auto-decoder under both strategies and present the results in Fig. 6.
The first row shows the results using the points-set sampling strategy, while the second row corresponds to directly identifying the nearest points on the SMPL mesh. The normal maps in the second row exhibit noticeable holes, indicating insufficient sampling. These artifacts are particularly pronounced for loose clothing (middle case in Fig. 6) but also appear in regions of relatively tight clothing (left and right case in Fig. 6).
5.1.2 Geometry Distribution Formulations
Figure 4: Single-geometry fitting visualization of different distribution formulations at the $10,000$th step.
To evaluate the training efficiency of different geometry distributions, we first conduct a single-geometry fitting task and then extend our analysis to a dataset. We compare our method with several alternatives, including the original geometry distribution proposed by Zhang et al. [52] (Original), the naive formulation in Eq. 2 (w/o pairs), and the formulation sampling from $\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma})$, proposed in Eq. 3.
For the single-geometry fitting task, we train each formulation for $10,000$ iterations and visualize the synthesized point clouds at the final step in Fig. 4. Our method produces the smoothest and most accurate reconstructions, while all other methods generate points that deviate from the surface. The Chamfer distances at the final step are presented in Tab. 2 (Single). The original geometry distribution performs the worst because it samples from Gaussian noise and requires more iterations to capture the coarse human shape. The “w/o pairs” formulation also exhibits suboptimal performance as it must learn mappings between distant points, leading to slower convergence. The approach that samples from $\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma})$ achieves the lowest Chamfer distance in this task. With a fixed pose, each sample can focus on learning specified local features, thereby simplifying the optimization. However, as shown in Fig. 4, its reliance on dispersed Gaussian centers leads to noisy output for some regions. Furthermore, this approach negatively impacts convergence when scaling to datasets with diverse poses, as reflected in its degraded performance in Tab. 2 (Dataset).
For dataset-scale evaluation, we train each model on the THuman2 [49] dataset for $30,000$ iterations. As shown in Fig. 5, our method consistently achieves significantly higher fidelity compared to the other approaches within the same number of training iterations. The evolution of Chamfer distance over iterations is provided in the Appendix.
Original w/o pairs $\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma})$ Ours
Single 0.0083 0.0040 0.0020 0.0032
Dataset 0.0101 0.0706 0.0071 0.0032
Table 2: Chamfer distance of different distribution formulations.
Figure 5: Visualization results of different geometry distribution formulations at the $30,000$th training iteration. GT represents the ground-truth result. Notably, the results do not represent the final outcome, as training is not fully completed.
5.1.3 Network Architecture
We experiment to validate that the auto-decoder outperforms the existing auto-encoders for our task. The auto-encoder learns features from the input, requiring more computations and being potentially constrained by the input’s representation capacity.
We adopt the 3D representation network VectorSet [50] as the encoder for the auto-encoder and take $50,000$ input points, balancing representation capability and memory consumption. Since VectorSet uses independent embeddings rather than a feature map, we treat each pixel of the feature map $\mathbb{R}^{24\times 24\times 8}$ as an individual embedding, yielding 576 embeddings (FeatureMap). To further ensure a fair comparison, we introduce an additional model where we replace our feature map decompression module (Fig. 11 (b)) with the one proposed by VectorSet, maintaining as much of the VectorSet architecture as possible (VectorSet).
The experiment is conducted on the THuman2 dataset [49]. As shown in Tab. 3, we report the average distance between the synthesized points and the surface. While converting independent embeddings to a feature map improves reconstruction accuracy, the results remain inferior to our proposed auto-decoder.
Model Surface Distance ↓
VectorSet 0.0018
FeatureMap 0.0014
Ours 0.0012
Table 3: Surface distance comparison between various designs.
Figure 6: Comparison of strategies for constructing training pairs. The first row shows the results of our proposed strategy, and the second row finds the nearest points on the SMPL mesh.
5.2 Pose-conditioned Random Generation
This experiment measures the geometry quality of pose-conditioned generation. Consistent with previous work [57, 56], we utilize the THuman2 Dataset [49] for training and evaluation. We calculate the Fréchet Inception Distance (FID) metrics between the rendered normal images and ground truth normal maps. Notably, some works [41, 57, 56] enhance their rendering results using information such as normal field/map or predicted points’ rotations. Since our goal is to compare geometry quality, we primarily evaluate these methods based on their raw geometry outputs. To ensure fairness, we also report their performance on their final rendered results. For each subject, we render $50$ images from different views and synthesize $25,000$ normal images for the dataset.
We compare our method with representative human generation methods employing different 3D representations, including Gaussian splatting (E3Gen [56]), implicit function (GetAvatar [8], gDNA[41]), and Nerf (ENARF [24], GNARF [3] and Eva3D [11]). As shown in Tab. 4, our method significantly outperforms other approaches in terms of raw geometry outputs (blue background metric). In comparison to the state-of-the-art (gDNA), our method improves geometry quality by $57\%$ ($42.9$ to $16.2$). Additionally, when compared to the final rendered results from other methods, our raw geometry shows improvements as well ($7\%$ from $17.4$ to $16.2$). As shown in Fig. 7, E3Gen [56] synthesizes unnatural shapes with inconsistent normal colors. GetAvatar [57] generates cloth wrinkles with unnatural directional patterns, while gDNA [41]’s normals produce wrinkles that appear random, unrealistic, and not coordinated with the human pose. In contrast, our method generates realistic and pose-consistent clothing details, demonstrating superior geometric fidelity and detail preservation.
Method FID (normal) $\downarrow$
ENARF* 223.72
GNARF* 166.62
EVA3D* 60.37
E3Gen (raw geometry) 65.32
E3Gen (enhanced rendering) 28.12
GetAvatar (raw geometry) 56.07
GetAvatar (enhanced rendering) 22.77
gDNA (raw geometry) 42.90
gDNA (enhanced rendering) 17.43
Ours 16.16
Table 4: Comparison of FID scores on THuman2 dataset. The * results are adopted from E3Gen [56]. Metrics with a blue background are computed from the rendered results of the generated geometries without additional rendering enhancements.
Figure 7: Qualitative comparisons of pose-conditioned human generation results. The left two columns display final rendering results for baseline methods, while the right two columns show their raw geometry outputs. The last row illustrates two samples generated by our method.
5.3 Novel Pose Generation
Previous methods [57, 41, 56] synthesize 3D humans in canonical space and obtain novel pose appearances via deformation. As a result, while pose transformation is inherently stable, these methods lack the ability to generate pose-dependent clothing deformations (e.g., wrinkles and folds), as their geometry details remain unchanged after deformation. See the Appendix for visualization.
Unlike these methods, our approach directly synthesizes points on the deformed shape without relying on rigging mechanisms to ensure plausibility under pose changes. However, when the feature map is fixed and only the pose is altered, our method still produces reasonable results, as shown in Fig. 8 (Fixed Feature), demonstrating its robustness in handling novel poses. Furthermore, when provided with a single-view normal map specifying the clothing type, our network effectively learns a pose-aware feature map, enabling the synthesis of realistic clothing deformations, as shown in Fig. 8.
Figure 8: Comparison of novel pose generation results: (left) the source identity; (middle) novel poses generated with the source identity’s feature maps; (right) novel poses generated with the synthesized feature map from a single-view image.
In Fig. 9, we present an extreme case of novel pose generation, where the body heavily occludes most clothing, the target poses differ significantly from the input, and the clothing drapes far from the body. More results can be seen in the Appendix.
Figure 9: Extreme case of novel pose generation.
6 Limitations and Future Work
Similar to other generative methods, our approach is constrained by the diversity of the training datasets. Additionally, we cannot guarantee uniform sampling on the target geometry surface. This is because the number of target points $\mathbf{x}_{1}$ associated with each SMPL point $\mathbf{x}_{0}^{\prime}$ is not uniform. Some $\mathbf{x}_{0}^{\prime}$ may correspond to more $\mathbf{x}_{1}$ points than others. While sufficient sampling and discarding points that are too close to each other alleviate this problem, we encourage future work to explore more advanced strategies for constructing training pairs.
7 Conclusions
In conclusion, we introduce the generative human geometry distribution, the first method that learns the distribution of geometry distributions, enabling ‘infinite’ sampling to synthesize high-fidelity human geometries with realistic, pose-aware clothing details. Using flow matching, we learn a flow that transforms from the SMPL template to the target geometry, significantly improving training efficiency. Following this, our method encodes geometry into compact feature maps, facilitating various downstream generative tasks. Experimental results demonstrate that our approach achieves state-of-the-art quantitative performance in the pose-conditioned generation task compared to existing generative methods. Furthermore, in novel pose generation experiments, our method successfully synthesizes fine-grained clothing details that conform to human poses and exhibits strong robustness, generating plausible results even when provided with a feature map mismatched to the target pose. These results highlight the effectiveness of our representation and its potential for advancing 3D human modeling and synthesis.