### v3

# Generative Human Geometry Distribution

Abstract
Realistic human geometry generation is an important yet challenging task, requiring both the preservation of fine clothing details and the accurate modeling of clothing-body interactions. To tackle this challenge, we build upon Geometry distributions—a recently proposed representation that can model a single human geometry with high fidelity using a flow matching model. However, extending a single-geometry distribution to a dataset is non-trivial and inefficient for large-scale learning. To address this, we propose a new geometry distribution model by two key techniques: (1) encoding distributions as 2D feature maps rather than network parameters, and (2) using SMPL models as the domain instead of Gaussian and refining the associated flow velocity field. We then design a generative framework adopting a two-staged training paradigm analogous to state-of-the-art image and 3D generative models. In the first stage, we compress geometry distributions into a latent space using a diffusion flow model; the second stage trains another flow model on this latent space. We validate our approach on two key tasks: pose-conditioned random avatar generation and avatar-consistent novel pose synthesis. Experimental results demonstrate that our method outperforms existing state-of-the-art methods, achieving a $57\%$ improvement in geometry quality.
1 Introduction
3D human geometry generation is a critical yet challenging task. The human body exhibits high-frequency clothing details, which are inherently difficult to synthesize. Therefore, the core of this task is the design of a representation that can capture both the underlying structure and fine details, posing two key challenges: 1) encoding high-frequency details into a low-dimensional manifold without losing fidelity; 2) modeling the relationship between clothing wrinkles and poses to preserve realistic details.
Figure 1: Geometry distribution.
Existing methods represent the human body in different ways. NeRFs Men et al. (2024); Zhang et al. (2022); Xiong et al. (2023); Jiang et al. (2023); Xu et al. (2023); Zheng et al. (2024) focus on the rendering results but neglect underlying geometry and are constrained by the rendering speed and resolution. Implicit functions Xu et al. (2022), such as signed distance functions, struggle to synthesize thin structures and tend to oversmooth results. Point clouds Zhang et al. (2024c; d) and volume-based representations trade off between memory efficiency and quality. Recently, geometry distributions Zhang et al. (2025) model single 3D shapes as distributions of points on their surfaces. Samples $\mathbf{X}$ from a Gaussian distribution $\mathcal{N}$ are transformed via a flow-based diffusion network $\mathcal{E}(\mathbf{X})$ into the target geometry, capturing fine-grained surface details (see Fig. 1). This formulation enables infinite point sampling, allowing for a high-fidelity representation of individual shapes. However, extending a single-geometry distribution to an entire dataset is non-trivial because of two reasons: 1) single-geometry formulations store the geometry in the parameters of a flow network, which leads to prohibitive memory consumption and limits their scalability for generative tasks, and 2) while learning the flow velocity fields from Gaussian distributions to a single shape is feasible, extending this to multiple shapes across a dataset becomes computationally expensive and inefficient.
Figure 2: (a) Denoising process of a human geometry distribution. (b) Random avatar generation for a given pose. (c) Novel pose generation of a given avatar. Results are rendered from point clouds.
To this end, we propose a novel human geometry distributions model by two key techniques:
  • We encode each human geometry distribution into a feature map rather than the network weights of a flow network directly, providing a generalized way to represent geometry distributions.
  • We adopt the SMPL Pavlakos et al. (2019) model distribution instead of a Gaussian and refine the associated flow velocity field for efficient learning. Specifically, we construct spatially efficient mappings from the velocity field and reformulate them into a regularized dense space.
Based on these designs, we propose the first 3D generative method for geometry distributions, modeling the distribution of individual human geometry distributions. Our framework is formulated in two stages: first, we use a flow model to compress each human geometry distribution into a compact feature map, from which a high-fidelity human geometry can be sampled through a denoising process (Fig. 2 (a)), and then, we train another flow model to generate human geometry distributions (i.e., the feature maps). While the currently top-performing methods Xu et al. (2022); Zhang et al. (2024d; 2023b) obtain visually plausible geometric details by enhanced rendering techniques, our method synthesizes high-fidelity geometry directly, leading to large improvements. We design two generative tasks to evaluate the effectiveness of our method (Fig. 2 (b) and (c)). The first task generates random 3D human geometries conditioned on a given pose, while the second enables novel pose generation of a given avatar. Our quantitative results show that our method improves the geometry quality by $57\%$ ($42.9$ to $16.2$) and the visual appearance by $7\%$ ($17.4$ to $16.2$) compared to the state-of-the-art.
2 Related work
2.1 General 3D Generation
3D generation methods employ various representations, such as meshes, voxels, point clouds, and implicit functions, each offering distinct advantages but also facing inherent limitations Zhang et al. (2025). Here, we focus on compact representations that scale efficiently to large datasets.
Given one or more images, direct inference of pixel-aligned positions Yu et al. (2021a); Tang et al. (2024); Xu et al. (2024b; a) avoids intermediate representations, which is unsuitable for modeling intrinsic human shapes. A popular alternative is to synthesize tri-planes as intermediate representations Wang et al. (2023); Li et al. (2023); Hong et al. (2023), from which implicit functions or points Zou et al. (2024); Han et al. (2024) can be sampled. However, tri-planes are limited by their resolution, restricting their ability to capture fine-grained details. GaussianCube Zhang et al. (2024a) transfers Gaussians into a regular voxel grid via optimal transport, limiting its flexibility for intricate articulation. Gamba Shen et al. (2024) and VectorSet Zhang et al. (2023a) employ cross-attention mechanisms to inject 3D embeddings or image tokens, albeit constrained by memory consumption. Trellis Xiang et al. (2024) leverages sparse voxels but lacks mechanisms to incorporate human-specific priors. In contrast, 2D maps Rai et al. (2025); Elizarov et al. (2024); Yan et al. (2024), offer a memory-efficient alternative that aligns naturally with human priors.
2.2 3D Human Reconstruction and Generation
Features extracted from images Saito et al. (2019; 2020); Alldieck et al. (2022) provide rich view-dependent information, enabling high-fidelity human reconstruction. Some works incorporate additional cues such as voxel grids He et al. (2020), human templates Zheng et al. (2021), or multi-view images Shao et al. (2022) to mitigate the depth ambiguity of single-view image. For example, ICON Xiu et al. (2022) and ECON Xiu et al. (2023) predict the frontal and back normals from the SMPL template and reconstruct the human using these maps. Building on these reconstruction techniques, generative networks—such as GANs Xiong et al. (2023); Jiang et al. (2023) and diffusion models Sengupta et al. (2024); Shao et al. (2022); Zhang et al. (2024b)—have been employed to synthesize pixel-aligned features or generate multi-view images Kolotouros et al. (2024) for human generation. However, these methods are inherently limited by their reliance on view-dependent features, sensitivity to multi-view inconsistencies, and often require post-processing to fuse information from different views Zhang et al. (2024b), failing to capture the intrinsic 3D shape.
To learn intrinsic 3D representations, a popular approach Noguchi et al. (2022); Zhang et al. (2022); Yang et al. (2024); Wu et al. (2023; 2024) is to train a GAN model that synthesizes tri-planes and uses NeRF for rendering. However, these methods are constrained by rendering speed and resolution, often requiring super-resolution modules Bergman et al. (2022); Dong et al. (2023), multi-part structures Xu et al. (2023), or refinement stages Men et al. (2024); Zheng et al. (2024) to achieve high-fidelity results. Rather than employing tri-planes, alternative methods incorporate human templates, alleviating sparse spatial occupancy, but struggle with loose clothing Hong et al. (2022) or still rely on super-resolution modules Hu et al. (2024). Explicit representations, such as primitive volumes Chen et al. (2023), provide efficient rendering but face challenges in extracting detailed geometry. Mesh-based methods, using displacement Sanyal et al. (2024) or layered surface volumes Xu et al. (2024c), are limited by fidelity and struggle with loose clothing. Point-based representations Abdal et al. (2024); Zhang et al. (2024d; c) offer flexibility for modeling thin structures but are constrained by point density. Furthermore, it can be difficult to produce accurate 3D geometry under indirect supervision from rendering. In contrast, gDNA Xu et al. (2022) proposes learning implicit functions directly from 3D data, but synthesizing high-fidelity geometry remains challenging due to inherent representation limitations Zhang et al. (2025). Our approach directly learns from 3D data while enabling “infinite” sampling, leading to high-fidelity geometry synthesis. See Tab. 1 for comparison.
Table 1: Comparison of 3D human representation methods.
Method Loose Clothing Scalability Fine Details
Tri-plane + NeRF Noguchi et al. (2022); Zhang et al. (2022)
Template-based Hong et al. (2022); Hu et al. (2024)
Primitive Volumes Chen et al. (2023)
Mesh-based Sanyal et al. (2024); Xu et al. (2024c); Feng et al. (2023)
Point-based Abdal et al. (2024); Zhang et al. (2024d)
Implicit Function Xu et al. (2022); Zhang et al. (2023b)
Geometry Distributions (Ours)
3 Preliminaries
Flow Matching
Flow matching Lipman et al. (2023) is a variant of diffusion models. It constructs a flow that transforms samples from a source distribution $\mathbf{p}$ to a target distribution $\mathbf{q}$. Specifically, the training objective loss is defined as follow:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{x}_{0}\sim\mathbf{p},\mathbf{x}_{1}\sim\mathbf{q},t\in[0,1]}\left\|u_{\theta}(\mathbf{x}_{t},t)-(\mathbf{x}_{1}-\mathbf{x}_{0})\right\|$ (1)
where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+\mathbf{x}_{1}$. After training, an ordinary differential equation is solved to transition from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$, allowing us to sample $\mathbf{x}_{1}\sim\mathbf{q}$ by first sampling $\mathbf{x}_{0}\sim\mathbf{p}$.
Based on the flow matching formulation, the core of our modeling is to identify a suitable source distribution and a target distribution that aligns well with our problem. For brevity, we omit $t\in[0,1]$ in subsequent equations.
Geometry Distributions
Geometry distributions model a surface $\mathcal{M}\subset\mathbb{R}^{3}$ as a probability distribution $\Phi_{\mathcal{M}}$, where any sample $\mathbf{x}_{1}\sim\Phi_{\mathcal{M}}$ corresponds to a point on the surface. To achieve this,  Zhang et al. (2025) adapt a diffusion model to learn the transformation from a Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{1})$ to the target distribution of surface points. A point $\mathbf{x}_{1}\in\mathcal{M}$ on the target surface can then be obtained by solving an ODE with an initial point $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
4 Method
4.1 Overview
In this section, we first introduce our human geometry distribution formulation by identifying suitable source and target distributions (Sec. 4.2). Next, we describe encoding this distribution into a feature map using an auto-decoder architecture, as depicted in Fig. 4 (a) (Sec. 4.3). Finally, we present generative tasks aimed at synthesizing the feature map, guided by the SMPL template, optionally incorporating additional conditioning inputs, as illustrated in Fig. 4 (b) (Sec. 4.4).
4.2 Human Geometry Distribution
Our insight is to replace the Gaussian distribution $\mathcal{N}(0,1)$ in the prior geometry distributions Zhang et al. (2025) with the SMPL template shape distribution $\Phi_{\mathcal{S}}$, aligning the source distribution closer to the target geometry distribution $\Phi_{\mathcal{T}}$. A naive approach can be formulated as:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{x}_{0}\sim\Phi_{\mathcal{S}},\mathbf{x}_{1}\sim\Phi_{\mathcal{T}}}\left\|u_{\theta}(\mathbf{x}_{t},t)-(\mathbf{x}_{1}-\mathbf{x}_{0})\right\|,$ (2)
where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+\mathbf{x}_{1}$. Building upon this, we propose two strategies to reduce the modeling complexity and improve training efficiency, including explicitly constructing training pairs and normalizing samples from distributions.
Training Pair Construction.
Since the learned probability flow by Eq. 2 approximates a conditional optimal transport Lipman et al. (2024), one strategy for improving training efficiency is avoiding learning extraneous paths between distant points. To achieve this, we prioritize short flow by constructing training pairs set $\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}$ where $\mathbf{x}^{\prime}_{0}\sim\Phi_{\mathcal{S}}$ is geometrically close to $\mathbf{x}_{1}\sim\Phi_{\mathcal{T}}$. Specifically, we first sample a set of points $\{\mathbf{x}_{0}\}_{\mathcal{S}}$ on the SMPL template and a set of points $\{\mathbf{x}_{1}\}_{\mathcal{T}}$ on the target geometry. Then, for each $\mathbf{x}_{1}\in\{\mathbf{x}_{1}\}_{\mathcal{T}}$, the corresponding $\mathbf{x}_{0}^{\prime}$ is obtained by:
[…]
Notably, because multiple points $\mathbf{x}_{1}$ may share the same nearest SMPL points $\mathbf{x}_{0}^{\prime}$, especially in loose or wrinkled regions, directly learning a path from $\mathbf{x}_{0}^{\prime}$ to $\mathbf{x}_{1}$ leads to under-sampled geometry (similar to the artifacts shown in the second row of Fig. 7). Therefore, we add a perturbation drawn from $\mathcal{N}(0,\bm{\sigma})$ to $\mathbf{x}_{0}^{\prime}$, introducing randomness to enhance sample diversity. In this formulation, the source distribution becomes $\mathcal{N}(\mathbf{x}_{0}^{\prime},\bm{\sigma})$ and the target distribution is the human geometry. Sampling from the constructed training pairs, the flow matching optimization objective is denoted by:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{x}_{0}^{\prime},\bm{\sigma}),(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})\in\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}}\left\|\cdot\right\|$ (3)
Distribution Normalization.
Figure 3: Aggregated samples $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})$.
When training the transformation from the SMPL template to the human geometry of a dataset (Eq. 3), the network receives spatially imbalanced supervision due to unevenly sampling points, which arises from two factors. First, points lie only on the surfaces, making them extremely sparse and unevenly distributed relative to the full 3D space. Second, variations in human pose and body shape further exacerbate this spatial imbalance, as some regions are more dynamic and thus sampled less consistently across geometries. Fig. 3 visualizes this effect by aggregating sampled points from multiple training iterations, showing that some spatial regions have high point density while others remain sparsely sampled. This spatial imbalance leads to inefficient training, as the model pays more attention to high-density regions while neglecting low-density regions. To address this, we normalize both the source and target distribution by subtracting $\mathbf{x}_{0}^{\prime}$. The source distribution becomes a zero-centered Gaussian distribution $\mathcal{N}(0,1)$, with $\bm{\sigma}$ set to $1$, and the target distribution is modeled as a regularized dense displacement field $\Delta\mathbf{x}=\mathbf{x}_{1}-\mathbf{x}^{\prime}_{0}$. Note that this subtraction removes the positional information of $\mathbf{x}^{\prime}_{0}$. To retain this information without suffering from imbalanced sampling, we reintroduce $\mathbf{x}_{0}^{\prime}$ as a conditional signal that scales the network hidden feature, indirectly influencing the feature representation. By building flow between regularized dense spaces, we reduce the modeling complexity and achieve improved training efficiency, as demonstrated in Sec 5.2. The optimization objective is denoted by:
$\arg\min_{\theta}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}\left\|u_{\theta}(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime})-(\Delta\mathbf{x}-\mathbf{n})\right\|,\vskip-5.0pt$ (4)
where $\mathbf{n}\sim\mathcal{N}(0,1)$, $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})\in\{(\mathbf{x}^{\prime}_{0},\mathbf{x}_{1})\}_{\mathcal{T}}$ and $\mathbf{x}_{t}=(1-t)\mathbf{n}+t\Delta\mathbf{x}$.
Figure 4: Overview of our method. (a) We encode a geometry into a feature map, which is decompressed with a SMPL vertex map. The decompressed feature serves as a condition for our denoising network. (b) The human generation task is formulated as the conditional generation of feature maps, guided by the SMPL vertex map, optionally incorporating additional conditioning inputs.
4.3 Conditional Distribution Encoding
The dataset for our task consists of a set of pairs $\mathcal{D}=\{(\mathcal{S},\mathcal{T})\}$, each associated with a latent representation $\mathbf{z}_{\mathcal{T}|\mathcal{S}}$, which is used to recover $\mathcal{T}$ (target geometry) by conditioning on $\mathcal{S}$ (SMPL). This design aligns with our generative task, i.e. , $\mathcal{S}$-conditioned $\mathcal{T}$ generation. The training loss can be written as,
$\operatorname*{arg\,min}_{\theta,\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}}\mathbb{E}_{(\mathcal{S},\mathcal{T})\in\mathcal{D}}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}\left\|u_{\theta}(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime},\mathbf{z}_{\mathcal{T}|\mathcal{S}})-(\Delta\mathbf{x}-\mathbf{n})\right\|,$ (5)
To enhance the expressiveness, we model $\mathbf{z}_{\mathcal{T}|\mathcal{S}}\in\mathbb{R}^{C\times H\times W}$ as a compressed 2D feature map (height $H$ and width $W$), as shown in Fig. 4 (a). The representation aligns with recent advancements in UV representations for 3D objects Yan et al. (2024). Furthermore, we use a decoder network $\mathrm{Dec}_{\phi}(\cdot)$ to decompress $\mathbf{z}_{\mathcal{T}|\mathcal{S}}$ to a higher resolution. By leveraging the UV coordinates associated with $\mathbf{x}_{0}^{\prime}$, we can then obtain per-point latents on the surface $\mathcal{S}$ by sampling the high-resolution map, denoted as $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Thus, the revised optimization can be written as
$\operatorname*{arg\,min}_{\theta,\phi,\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}}\mathbb{E}_{(\mathcal{S},\mathcal{T})\in\mathcal{D}}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}\left\|u_{\theta}\left(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime},\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})\right)-(\Delta\mathbf{x}-\mathbf{n})\right\|.$ (6)
Decoder.
As illustrated in Fig. 4 (a), the network $\mathrm{Dec}_{\phi}(\cdot)$ is a UNet-style network that contains several downsampling and upsampling layers. We render SMPL vertex positions into UV maps and concatenate them with hidden features in the convolution layers.
Denoising network.
The design of the denoiser $u_{\theta}$ is adapted from Zhang et al. (2025). In this case, we have two additional conditioning signals, SMPL point $\mathbf{x}^{\prime}_{0}$ and latent $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Additionally, we augment $\mathbf{x}^{\prime}_{0}$ by concatenating normals and canonical coordinates. Surface normals provide directional cues for clothing inference, while canonical coordinates encode body part semantics (e.g., distinguishing limbs from the torso).
4.4 Generative Tasks
After finishing learning the latents $\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}$, we train generative models in the latent space. We investigate two tasks: pose-conditioned random generation and novel pose generation of a given avatar, both built on a U-Net Karras et al. (2024; 2025) architecture, as illustrated in Fig. 4 (b).
Pose-conditioned random generation synthesizes diverse human geometries conditioned on an SMPL template mesh. To encode pose information, we render SMPL vertex positions into UV maps and inject them as residual connections into the U-Net architecture. For novel pose generation, the model additionally takes a frontal normal image as an extra condition to indicate the avatar identity. Specifically, for each animation sequence in the dataset, we randomly select one frame to provide the normal image condition and another frame to provide the SMPL pose condition. We leverage the frozen DINO-ViT model Caron et al. (2021) to extract image features, which are then fused into the U-Net via cross-attention layers.
The pose-conditioned generation networks are trained on the THuman2 dataset Yu et al. (2021b), while the novel pose generation is trained on 4DDress Wang et al. (2023). See the Appendix for implementation details.
5 Experiments and Results
In this section, we first conduct ablation studies on the training pair construction strategy and network architecture. Next, we compare different formulations of geometry distribution, including the formulation proposed by  Zhang et al. (2025) and our proposed alternatives. Finally, we compare our method with state-of-the-art generative methods regarding the two selected generative tasks.
5.1 Ablation Study
5.1.1 Training Pairs Construction
Rather than searching for the nearest points on the dense SMPL mesh, we first sample a relatively sparse set of points from the SMPL mesh and then determine the nearest points $\mathbf{x}_{0}$’ from this sampled set. It distributes the mapping workload across different $\mathbf{x}_{0}$’, which improves sampling efficiency. To validate this, we train our auto-decoder under both strategies and present the results in Fig. 7.
The first row shows the results using the sparse points set sampling strategy, while the second row corresponds to directly identifying the nearest points on the SMPL mesh. The normal maps in the second row exhibit noticeable holes, indicating insufficient sampling. These artifacts are particularly pronounced for loose clothing (middle case in Fig. 7) but also appear in regions of relatively tight clothing (left and right case in Fig. 7).
5.1.2 Network Architecture
Figure 5: Single-geometry fitting visualization.
We experiment to validate that the auto-decoder outperforms the existing auto-encoders for our task. The auto-encoder learns features from the input, requiring more computations and being potentially constrained by the input’s representation capacity. We adopt the 3D representation network VecSet Zhang et al. (2023a) as the encoder for the auto-encoder and take $50,000$ input points, balancing representation capability and memory consumption. Since VecSet uses independent embeddings rather than a feature map, we treat each pixel of the feature map $\mathbb{R}^{24\times 24\times 8}$ as an individual embedding, yielding 576 embeddings (FeatureMap). To further ensure a fair comparison, we introduce an additional model where we replace our feature map decompression module (Fig. 4 (b)) with the one proposed by VecSet, maintaining as much of the VecSet architecture as possible (VecSet).
The experiment is conducted on the THuman2 dataset Yu et al. (2021b). As shown in Tab. 3, we report the average distance between the synthesized points and the surface. While converting independent embeddings to a feature map improves reconstruction accuracy, the results remain inferior to our proposed auto-decoder.
5.2 Comparisons of Geometry Distribution Formulations
To evaluate the training efficiency of different geometry distributions, we first conduct a single-geometry fitting task and then extend our analysis to a dataset. We compare our method with several alternatives, including the geometry distribution proposed by Zhang et al. (2025), the naive formulation in Eq. 2 (w/o Pairs), and the formulation sampling from $\mathcal{N}(\mathbf{x}_{0}^{\prime},\bm{\sigma})$ without distribution normalization (w/o DistNorm), proposed in Eq. 3.
For the single-geometry fitting task, we train each formulation for $10,000$ iterations and visualize the synthesized points at the final step in Fig. 5. Our method produces the smoothest and most accurate reconstructions, while all other methods generate points that deviate from the surface. The Chamfer distances at the final step are presented in Tab. 3 (Single). The distribution proposed by Zhang et al. (2025) performs the worst because it samples from Gaussian noise and requires more iterations to capture the coarse human shape. The “w/o Pairs” formulation also exhibits suboptimal performance as it must learn mappings between distant points, leading to slower convergence. The approach that samples from $\mathcal{N}(\mathbf{x}_{0}^{\prime},\bm{\sigma})$ (w/o DistNorm) achieves the lowest Chamfer distance in this task. With a fixed pose, each sample can focus on learning specified local features, thereby simplifying the optimization. However, as shown in Fig. 5, its reliance on dispersed Gaussian centers leads to noisy output for some regions. Furthermore, this approach negatively impacts convergence when scaling to datasets with diverse poses, as reflected in its degraded performance in Tab. 3 (Dataset).
For dataset-scale evaluation, we train each model on the THuman2 Yu et al. (2021b) dataset for $30,000$ iterations. Our method consistently achieves significantly higher fidelity compared to the other approaches within the same number of training iterations, as shown in Fig. 7. Notably, the results do not represent the final outcome, as training is not fully completed. The evolution of Chamfer distance over iterations is provided in the Appendix.
Figure 6: Comparison of strategies for constructing training pairs. The first row shows our proposed strategy, the second row finds the nearest points on the SMPL mesh.
Figure 7: Visualization results of different geometry distribution formulations at the $30,000$th training iteration. GT represents the ground-truth result.
Table 2: Chamfer distance of different distribution formulations.
Zhang et al. w/o Pairs w/o DistNorm Ours
Single 0.0083 0.0040 0.0020 0.0032
Dataset 0.0101 0.0706 0.0071 0.0032
Table 3: Surface distance comparison between various designs.
Model Surface Distance ↓
VecSet 0.0018
FeatureMap 0.0014
Ours 0.0012
5.3 Pose-conditioned Random Generation
This experiment measures the geometry quality of pose-conditioned random generation. Consistent with previous work Zhang et al. (2023b; 2024d), we utilize the THuman2 Dataset Yu et al. (2021b) for training and evaluation. We calculate the Fréchet Inception Distance (FID) metrics between the rendered normal images and ground truth normal maps. Notably, some works Xu et al. (2022); Zhang et al. (2023b; 2024d) enhance their rendering results using information such as normal field/map or predicted points’ rotations. Since our goal is to compare geometry quality, we primarily evaluate these methods based on their raw geometry outputs. To ensure fairness, we also report their performance on their enhanced rendered results. For each subject, we render $50$ images from different views and synthesize $25,000$ normal images for the dataset.
We compare our method with representative human generation methods employing different 3D representations, including Gaussian splatting (E3Gen Zhang et al. (2024d)), implicit function (GetAvatar Feng et al. (2023), gDNAXu et al. (2022)), and Nerf (ENARF Noguchi et al. (2022), GNARF Bergman et al. (2022) and Eva3D Hong et al. (2022)). As shown in Tab. 4, our method significantly outperforms other approaches in terms of raw geometry outputs. In comparison to the state-of-the-art (gDNA), our method improves geometry quality by $57\%$ ($42.9$ to $16.2$). Additionally, when compared to the enhanced rendered results from other methods, our raw geometry shows improvements as well ($7\%$ from $17.4$ to $16.2$). As shown in Fig. 8, E3Gen Zhang et al. (2024d) synthesizes unnatural shapes with inconsistent normal colors. GetAvatar Zhang et al. (2023b) generates cloth wrinkles with unnatural directional patterns, while gDNA Xu et al. (2022)’s normals produce wrinkles that appear random, unrealistic, and not coordinated with the human pose. In contrast, our method generates realistic and pose-consistent clothing details, demonstrating superior geometric fidelity and detail preservation.
Table 4: Comparison of FID scores. The * results are adopted from E3Gen Zhang et al. (2024d). For some methods, the results are rendered directly from their raw geometries, so the numbers in both rows are identical.
ENARF* GNARF* EVA3D* E3Gen GetAvatar gDNA Ours
Raw Geometry 223.72 166.62 60.37 65.32 56.07 42.90 16.16
Enhanced Rendering 223.72 166.62 60.37 28.12 22.77 17.43 16.16
Figure 8: Qualitative comparisons of human generation results. Existing methods are displayed with both rendering results and raw geometry outputs, typically organized as two columns each. We include two geometry samples generated by our method for comparison.
5.4 Novel Pose Generation
Table 5: User study on quality and physical plausibility.
Metric GetAvatar gDNA E3Gen Ours
Quality ↑ 2.16 2.54 2.12 4.04
Plausibility ↑ 2.20 2.66 2.08 4.36
Previous methods Zhang et al. (2023b); Xu et al. (2022); Zhang et al. (2024d) synthesize 3D humans in canonical space and deform them via rigging, which limits pose-dependent clothing details (e.g., wrinkles and folds). In contrast, our approach directly synthesizes points on the deformed human body given the pose-aware feature map, enabling the generation of pose-dependent clothing details, as shown in Fig. 9. More visual results can be found in the Appendix.
For quantitative comparisons, FID metrics reflecting geometric quality are provided in Tab. 4. Note that since prior methods’ geometric details remain static across poses, their FID results correspond closely to those reported in the above table. In terms of physical plausibility of geometric details, due to the absence of standard metrics, we conduct a user study to demonstrate that our method uniquely enables pose-aware deformations that appear physically reasonable. For each method, we generate 2 identities under 8 poses and ask 25 participants to rate geometry quality and physical plausibility (1–5 scale), as shown in Tab. 5.
Our superior physical plausibility mainly stems from our synthesized pose-aware feature maps. Nevertheless, even when using a feature map that is incompatible with the body pose, our method still generates visually reasonable results, which demonstrates the robustness of our model in handling novel poses (see the Appendix for details).
Figure 9: Generating different poses for a given avatar. We showcase challenging cases, including exaggerated poses, skirts, and loose outfits.
6 Limitations
A limitation of our method lies in non-uniform sampling on the target geometry surface, as the number of target points $\mathbf{x}_{1}$ associated with each SMPL point $\mathbf{x}_{0}$’ varies. Although sufficient sampling and removing points that are too close mitigate this issue, we encourage future work to develop more advanced strategies for training pair construction. Besides, similar to other generative methods, our approach is constrained by the diversity of the training datasets. Specifically, our model generalizes well to various body types, as the dataset includes a reasonable diversity in this aspect. However, it cannot generate clothing styles that are entirely absent from the training data. Furthermore, the use of UV maps can cause seam artifacts in some randomly generated results due to discontinuity segmentation. A better approach would use UV segmentation aligned with real-world garment patches. Since our focus is on a modeling representation, addressing this issue is left for future work.
7 Conclusions
In conclusion, we introduce the generative human geometry distribution, the first method that integrates geometry distributions into generative modeling. This distribution-over-distribution design is novel, unique, and particularly effective, making our approach the only one capable of producing fine-grained geometry in a generative framework. Specifically, using flow matching, we learn a flow that transforms from the SMPL template to the target geometry, significantly improving training efficiency. Following this, our method encodes geometry into compact feature maps, facilitating various downstream generative tasks. Experimental results demonstrate that our approach achieves state-of-the-art quantitative performance compared to existing generative methods. Furthermore, our method is able to synthesize fine-grained clothing details that conform to human poses and exhibits strong robustness, generating plausible results even when provided with a feature map mismatched to the target pose. These results highlight the effectiveness of our representation and its potential for advancing 3D human modeling and synthesis.
Appendix
Figure 10: Details of our auto-decoder. (a) We sample noise from a Gaussian distribution ($\mathbb{R}^{3}$) and generate the flow that transforms from the SMPL template distribution to the human geometry distribution. (b) A convolutional network is used to interpret and decompress the compact feature map with the SMPL vertex map. (c) The architecture details of our model.
Implementation details
Our training requires an infinite set of training pairs $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})$. In practice, we sample $2^{18}$ surface points $\mathbf{x}_{1}$ per human geometry. We observe that many points in loose clothing regions share the same nearest SMPL template location, requiring multiple samples from that location to reconstruct these regions. To alleviate this, we sample a slightly sparser set of $2^{17}$ points on the SMPL template and apply the k-nearest neighbors (KNN) algorithm to associate each $\mathbf{x}_{1}$ with the nearest template point. This distributes the mapping workload across different template locations, improving the sampling efficiency for loose clothing regions. During training, we randomly sample $30,000$ pairs per iteration.
As shown in Fig. 10, the compact feature map $z_{\mathcal{T}|\mathcal{S}}$ is set to $\mathbb{R}^{8\times 24\times 24}$. We adopt the linear and convolutional layers from EDM2 Karras et al. (2024; 2025), using $256$ channels in all layers. For optimization, we use AdamW Loshchilov (2017) with a learning rate of $8\times 10^{-3}$, following a cosine learning rate scheduler across all models. We apply L2 regularization with weight $1e^{-4}$ to the feature map during training.
Figure 11: The evolution of the Chamfer distance of training different distribution formulations.
We train both the auto-decoder and the generation model on the THuman2 dataset Yu et al. (2021b) for the pose-conditioned generation task, following previous methods Xu et al. (2022); Zhang et al. (2024d) to ensure fair comparisons. The auto-decoder model is trained for two days and the generation model for one day on a machine with two A100 GPUs. To enhance the model’s ability to generalize across diverse poses, we employ the classifier-guidance that randomly drops the pose condition with a probability of $0.5$. This encourages the network to learn the underlying data distribution rather than overfitting to specific poses. For the novel pose generation task, we utilize the 4DDress dataset Wang et al. (2024), which comprises 64 outfits across $520$ human motion sequences, totaling approximately $50,000$ human geometries with diverse poses and clothing styles. We train the auto-decoder on a machine with $4$ A100 GPUs for two days and train the novel pose generation model for one day. We don’t utilize the classifier-guidance strategy on this dataset as its larger scale naturally supports robust training.
The auto-decoder model has 56.0256.02M parameters. On an A100 GPU, the feature map decompression takes $81$ ms, which is executed only once for the entire diffusion process. Each denoising step requires $529$ ms for $1$ million points, with a peak memory usage of $9522$ MB. On memory-constrained devices, the model can sample multiple times with significantly fewer points per iteration. For instance, sampling $10,000$ points takes $71$ ms with a peak memory consumption of $1420$ MB.
We visualize our results by sampling $1$ million points for Gaussian Splatting (GS) Kerbl et al. (2023) rendering. Specifically, we render GS’s depth maps first, from which we then extract normal maps. All points share the same scaling value, determined by the average minimum distance between points.
Figure 12: Demonstrations of different methods with fixed geometry features across varying poses. All results show identical geometric details in different poses.
Figure 13: Novel pose generation results.
Geometry Distribution Formulation
For the experiment of different formulations in training on a dataset, the evolution of the Chamfer distance across $30,000$ iterations is illustrated in Fig. 11. Our proposed human geometry distribution achieves superior reconstruction quality, with a consistent decrease in the Chamfer distance.
Novel Pose Generation
Previous methods Zhang et al. (2023b); Xu et al. (2022); Zhang et al. (2024d) synthesize 3D humans in canonical space and obtain novel pose appearances via deformation. Their limitations are illustrated in Fig. 12, where we deform avatars to different poses, resulting in identical clothing details across all poses. In the main paper, we have showcased that our method can generate pose-aware deformations by synthesizing a pose-compatible feature map. Fig. 12 demonstrates that, although the wrinkle details may not be entirely natural, our method still produces reasonable results even when using an incompatible feature map.
Furthermore, we present additional novel pose generation results in Fig. 13. Specifically, we showcase challenging cases featuring intricate poses and complex clothing deformations, such as loose garments and skirts, which significantly deviate from the SMPL surface.
Use of Large Language Models
During the preparation of this manuscript, we employed a large language model (LLM) as an auxiliary tool to assist with grammar checking and minor language refinements. All scientific content, ideas, and claims were conceived and written by the authors. The final version of the manuscript was carefully reviewed, modified, and approved by the authors to ensure accuracy and clarity.